# Enhancing Efficiency in Multidevice Federated Learning through Data Selection

### v4

@@ 6.1. Model Accuracy
Centaur outperforms AP training and UCD training for different encoders. We permute the four encoders and three classifiers and report the test accuracy of Centaur compared to the two other baselines: AP training and UCD training. In Figure 6 we report the highest test accuracy for each encoder across all three classifiers for three datasets. Results show that Centaur outperforms AP training by $0.53\%\sim 40.15\%$ and UCD training by $0.45\%\sim 17.78\%$, depending on the settings. There are several settings in which all FL training methods cannot reach a good accuracy, e.g., NASNet and ShuffleNet on CIFAR100, probably because of their relatively small model sizes compared to data complexity. Among them, Centaur still achieves better performance.
Centaur outperforms UCD training for different classifiers. In Figure 7 we present the accuracy of different classifiers for each encoder. In the top plot, the results of training on CIFAR10 and CIFAR100 with an accuracy higher than $30\%$ are presented for better visibility. In the bottom plot, we report the results of UCIHAR, MotionSense, and PAMAP2. The results show that Centaur outperforms both UCD training in all classifiers’ sizes. The test accuracy also tends to be similar across small, medium, and large classifiers. In addition, the classifier’s sizes may have less impact on more sophisticated encoders (e.g., EfficientNet and MobileNet), which is also observed in a previous work (Mo et al., 2021). This may be because an appropriate encoder already produces high-quality features for unseen data that are easy to learn (e.g., CIFAR10), and in such a case, classifier sizes do not make any difference in test accuracy. Similar patterns are observed in HAR datasets. Notice that for HAR datasets, there is currently a lack of publicly accessible pre-trained encoders to start FL with. As a result, we can observe a more significant performance gap compared to image classification, particularly for more complex datasets, e.g., PAMAP2. This lack of pre-trained encoders can also be the reason that, in HAR datasets, the performance of Centaur and AP are almost the same.
@@ 6.2. Efficiency
Cost-Accuracy Trade-off. We use CIFAR10, and fix MobileNet-v3 as the encoder and medium size for the classifier. We compute the test accuracy for a range of MAC, Bandwidth, Latency, and Energy budgets. The average size of each sample of CIFAR10 is $30KB$. As shown in Figure 8 (top), on both UCD and AP and for all ranges of accuracy, Centaur achieves lower MAC and Latency than both UCD training and AP training. In addition, Figure 8 (bottom) shows that communication Bandwidth and Latency of Centaur is almost the same as AP training; however Centaur can achieve a higher test accuracy along with more training steps while consuming more resources. Centaur causes much more communication cost than UCD training, because the training only involves the transmission of the classifier, which is much lighter than the encoder and data samples transmitted by Centaur. However, the accuracy of UCD training cannot go further than $83.10\%$, while Centaur can reach up to $89.90\%$ accuracy. We remark that model training consumes much more energy and causes much more latency than communication.
Correlation among Cost Metrics. To show the connection between cost metrics, we plot MAC, Latency, and Energy in Figure 9. We observe that both Latency and Energy have linear relationships with MAC. Besides, Latency is also linearly correlated with Energy, in such a way that these two entirely overlap when we appropriately scale the y-axes. We remark that Bandwidth also has a similar correlation. Such linear relationships are because of the assumption we made when computing Latency and Energy; however, the actual cost and the relationship among these metrics might not deviate much in practice.
Based on the above-mentioned correlation, we can add Latency (or Energy) of model training as well as the communication Bandwidth to get the overall workload. The results shown in Figure 1 indicate that the workload still follows a similar pattern as the computation in model training, because the training workload significantly overweight the communication workload.
@@ 6.3. Data and Participation Heterogeneity
Centaur is robust to data heterogeneity. To create imbalanced non-IID data partitions among clients, we use LDA as defined in § 4. Next, we set different values for LDA-Alpha to manipulate the levels of non-IID data partitions. Figure 11 and Figure 11 show the results when LDA-Alpha are changing from $0.001$, to $0.1$, to $1$, and to $1000$. The smaller is LDA-Alpha, the less balanced is the dataset. We also annotate some examples of class distribution in Figure 11 demonstrating that LDA-Alpha$=0.001$ generates almost $1$ class per client, while LDA-Alpha$=1000$ generates almost uniform classes per client. Results show that in all cases Centaur can reach higher test accuracy than both AP and UCD training. We remark that due to the retraining of the encoder, both Centaur and AP training can cause more fluctuations than UCD training; this is especially obvious when data distribution tends to be IID (e.g., when LDA-Alpha$=1000$ in Figure 11).
Centaur scales well with higher total and participating clients. We further scale the FL training problem with the number of total clients from 10 to 1000. $10\%$, $20\%$, and $50\%$ of them are selected respectively as participating clients in each round. Figure 12 shows Centaur achieves higher test accuracy compared to both UCD training and AP training. Specifically, with a higher participation rate, all training methods tend to have better performance. However, with a much larger number of total clients (e.g., 1000), the test accuracy is reduced, because the same size of data is partitioned into more portions. Also, with such more fragmented data partitioning, Centaur has much accuracy gain when compared to UCD training.
Figure 10. The impacts of imbalanced data partitioning on the test accuracy of AP training, UCD training, and Centaur, reported on MobileNetv3 on CIFAR10 dataset (annotated texts give examples of the classes present on the UCDs)
Figure 10. The impacts of imbalanced data partitioning on the test accuracy of AP training, UCD training, and Centaur, reported on MobileNetv3 on CIFAR10 dataset (annotated texts give examples of the classes present on the UCDs)
Figure 11. Impacts of imbalanced data partitioning on the test accuracy of AP training, UCD training, and Centaur, reported on ConvNet and UCIHAR dataset dataset.
@@ 6.4. Performance of Centaur under Mobility
We evaluate situations when the connection probability $\lambda$ changes based on the mobility model defined in §4. For a more readable visualization, we bucket $\lambda$ in three different ranges of $\lambda\in$ [0.1,0.4], $\lambda\in$ [0.4,0.7], and $\lambda\in$ [0.7,1.0] and report the corresponding Accuracy of Centaur when the UCDs are mobile. Results reported in Figure 13 show that Centaur always has higher efficiency than standard UCD training while incurring a higher latency than direct AP training. Specifically, regarding the much important UCD side, Centaur gains $6.45\%$, $5.64\%$, and $4.11\%$ higher accuracy with $51.73\%$, $60.48\%$, and $68.35\%$ lower cost both in terms of energy and latency across the different ranges of $\lambda$.
Interestingly, we also observe a rise in accuracy with lower connectivity probability. The primary reason is attributed to the overall design of our experimental setup whereby we make a realistic assumption that when mobile with limited connectivity, these UCDs can gather more data and perform the data selection. Additionally, with this judiciously selected data, the device trains the classifier for an extended number of epochs thus boosting its accuracy.
7. Discussion and Related Work
Multitier FL. Compared to the default client-server FL, client-edge-server FL (Liu et al., 2020; Abad et al., 2020) considers a middle layer, e.g., a cellular base station, to orchestrate FL training across clients and a central server. In such a setup, clients in the last tier participate with several devices (Cho et al., 2022; Diao et al., 2020; Zhang et al., 2021). This is shown to achieve better trade-offs between accuracy and bandwidth consumption. However, these works assume local training of the whole model on the client devices and do not utilize the capabilities of a resourceful edge device for aiding a constrained device during training, as they are only used for aggregation/communication. This makes it impractical in scenarios involving UCDs.
On-device Training. Most prior works only consider the challenges of on-device inference (i.e., forward pass) (Fedorov et al., 2019; Banbury et al., 2021; Xu et al., 2022; Liberis et al., 2021), or work on sparse training (Li et al., 2021a; Bibikar et al., 2022). Sparse training or pruning still requires computational resources (especially memory) that UCDs do not have. Few works propose solutions for full on-device training (i.e., forward and backward pass) (Lin et al., 2022), but not for FL scenarios as they often make impractical assumptions like the availability of large datasets for network architecture search (Liberis et al., 2021) or unconstrained memory for hosting the full model (Diao et al., 2020; Zhang et al., 2021). Transfer learning can be used on UCDs to fix the encoder architecture, which in turn restricts the extraction of newer features, and only trains the last layer for the personalized training (Zhang et al., 2021). Regarding complex tasks, training the fully functional model on UCDs is challenging, because the on-device model still needs the capability of learning all features by itself (Li et al., 2021a; Bibikar et al., 2022), and the computational resources at the upstream devices are unutilized.
Figure 12. The impact of participating clients and total clients on test accuracy reported on MobileNetv3 and CIFAR10.
Communication/Power Efficiency. Using a pre-trained encoder without further training might fail to capture discerning features from the newly generated data, causing lower performance for the ultimate classifiers. On the other hand, passing all collected sensor data from UCDs to APs (even with data compression) is not always feasible and it might consume a lot of energy and bandwidth. Importantly, network availability for UCDs can be fragile due to the low-energy wireless communication protocols they use and the mobility of these UCDs. As UCDs are becoming stand-alone, it is expected that they can operate as independent devices without internet or permanent connection to an upstream device, Centaur takes a major step in unlocking more functionalities of such development. Compression techniques (Lim et al., 2019) or topology-aware hierarchical aggregation (Luo et al., 2020) can reduce Latency by taking advantage of the locality in networks and continuously balancing workloads across clients. However, compression and hierarchical aggregation alone do not allow for seamlessly incorporating UCDs in FL. Our proposed method comprises different methods to offer customizable trade-offs between test accuracy and communication cost in data- and topology-agnostic manners, while allowing additional use of the recent compression and hierarchical aggregation techniques for further efficiency.
Training on UCDs. By default, FL algorithms are not specifically designed for a multi-device environment (Lee et al., 2019; Islam and Nirjon, 2019). Training DNNs on tiny MCU devices is extremely challenging due to limited memory, computation, and power resources. There are recent efforts to enable running ML models on a micro-controller with highly limited memories, such as TensorFlow Lite Micro (David et al., 2021) or Tiny Training Engine (Lin et al., 2022). Such developments, usually based on a combination of hardware and software novelties, can be easily integrated into Centaur to provide further efficiency. As we observe, a better training framework can not only enable deploying larger classifiers on UCDs, but can also allow us to calculate more gradients for better data selection. While it is difficult to train large DNNs on UCDs, these devices are typically parts of a personal network, and therefore, they can privately communicate/coordinate with other surrounding devices. Our work considers such a hierarchy for unlocking more utilities while providing better efficiency.
@@ 7. Discussion and Related Work
8. Limitations and Future Work
First, we accommodate FedAvg (McMahan et al., 2017b) in our current implementation, but Centaur is compatible with other FL algorithms, e.g., FedProx (Li et al., 2020), FedAdam (Reddi et al., 2020), or DM-PFL (Zhang et al., 2023). On the other hand, layer-wise FL methods, such as FedMA (Wang et al., 2020), may have interesting interactions as they also conduct model training in a partitioned fashion.
Second, more algorithmic changes can be done to existing AP training, but we emphasize that whatever we do for AP training can only be useful for applications where all data is available at AP. In many emerging applications, UCDs are the main devices that collect and process data in the first place, and we do not want UCDs to send all the data to AP and put the burden of data selection on the APs shoulders.
Third, we do not perform model selection or architecture search, and our model partitioning simply follows the encoder-classifier style of benchmark networks, but one can enhance the model initialization and partitioning with state-of-the-art advancements in neural architecture search (Li et al., 2021b; Huang et al., 2023).
Fourth, our goal is to present a proof of concept for a multi-device FL that encompasses UCDs and APs. We present a real testbed and perform actual measurements, but in the rest of our experiments, we utilize GPU servers for running end-to-end FL scenarios. The constraints of UCDs are simulated according to the size and type of the model they can train. As a topic for future investigation, and to achieve more precise calculations of compute and storage resources, the end-to-end process of Centaur could be implemented on real-world UCDs and APs.
Fifth, we considered a one-to-one relationship between UCDs and APs, but one can expand Centaur into a many-to-one setting. This, however, needs tackling the challenge of aggregating diverse classifiers trained by various sensing devices.
Sixth, we evaluated Centaur on supervised FL tasks, but it is possible to suggest a supplementary approach for implementing semi-supervised or unsupervised tasks. Loss-based data selection has been studied very well in the ML community and shows promising advantages. Finally, dealing with outliers and distribution shifts are important challenges but we believe these are topics outside the scope of this paper.
@@ 9. Conclusions
To include constrained devices in an FL task, we offer an efficient training approach that utilizes data selection to enhance learning while minimizing computational costs. Additionally, we employ partition-based training and aggregation, wherein constrained devices train a specific part of the model, aided by more capable devices that train the complete model. Our assessments validate the benefits of our proposed FL framework, Centaur, in enhancing model accuracy while simultaneously reducing system overhead. Furthermore, we propose several promising directions for future studies in the area of efficient FL on constrained sensory devices.

### v6

@@ 6.2. Model Accuracy
We permute the four encoders and three classifiers and report the test accuracy of Centaur compared to the two other baselines: AP training and UCD training. In Figure 5 we report the highest test accuracy for each encoder across all three classifiers for three datasets. Results show that Centaur outperforms AP training by $0.53\%\sim 40.15\%$ and UCD training by $0.45\%\sim 17.78\%$, depending on the settings. There are several settings in which all FL training methods cannot reach a good accuracy, e.g., NASNet and ShuffleNet on CIFAR100, probably because of their relatively small model sizes compared to data complexity. Centaur still achieves better performance compared to the baselines.
In Figure 6 we present the accuracy of different classifiers for each encoder. In the top plot, the results of training on CIFAR10 and CIFAR100 with an accuracy higher than $30\%$ are presented for better visibility. In the bottom plot, we report the results of UCIHAR, MotionSense, and PAMAP2. The results show that Centaur outperforms both UCD training in all classifiers’ sizes. The test accuracy also tends to be similar across small, medium, and large classifiers. In addition, the classifier’s sizes may have less impact on more sophisticated encoders (e.g., EfficientNet and MobileNet), which is also observed in a previous work mo2021ppfl. This may be because an appropriate encoder already produces high-quality features for unseen data that are easy to learn (e.g., CIFAR10), and in such a case, classifier sizes do not make any difference in test accuracy. Similar patterns are observed in HAR datasets. Notice that for HAR datasets, there is currently a lack of publicly accessible pre-trained encoders to start FL with. As a result, we can observe a more significant performance gap compared to image classification, particularly for more complex datasets, e.g., PAMAP2. This lack of pre-trained encoders can also be the reason that, in HAR datasets, the performance of Centaur and AP are almost the same.
These results emphasize that with a properly pre-trained backbone, continual re-training of the encoder on APs is neither necessary nor desirable. In transfer learning, it is standard to fix, or only occasionally fine-tune, the encoder and update the lightweight classifier head to adapt to local data. In Centaur, we follow this principle: UCDs perform frequent classifier-only updates to rapidly personalize to local data, while APs periodically update the encoder to correct drift in feature extraction. This approach achieves competitive accuracy with lower communication and computation costs: frequent classifier updates capture personalization and mobility, while periodic encoder updates maintain convergence and generalization. Future work can explore how varying update frequencies affect convergence trade-offs.
@@ 6.3. Efficiency
Cost-Accuracy Trade-off. Here, with CIFAR10, MobileNet-v3 as the encoder, and a medium size for the classifier, we compute the test accuracy for a range of MAC, Bandwidth, Latency, and Energy budgets. The average size of each sample of CIFAR10 is $30KB$. Figure 7 (top) shows on both UCD and AP and across accuracy levels, Centaur achieves lower MAC and Latency than both UCD training and AP training. Figure 7 (bottom) shows that communication Bandwidth and Latency of Centaur is almost the same as AP training; however, Centaur can achieve a higher test accuracy along with more training steps when consuming more resources. Centaur uses much more communication resources than UCD training, as Centaur transmits both the classifier, although it is much lighter than the encoder, and data samples. However, the accuracy of UCD training cannot go further than $83.10\%$, while Centaur can reach up to $89.90\%$ accuracy. More reduction in uplink data transmission of UCDs can be achieved by a more advanced data selection strategy, lowering the communication burden on constrained devices. This also leads to a notable decrease in overall communication cost compared to AP training, as less data is exchanged between UCDs and APs while improving model accuracy. We remark that model training consumes much more energy and causes much more latency than communication.
Figure 9. The impacts of imbalanced data partitioning on test accuracy of AP training, UCD training, and Centaur, on MobileNetv3 with CIFAR10 (annotated texts are examples of the classes present on a UCD)
Figure 10. The impacts of imbalanced data partitioning on test accuracy of AP training, UCD training, and Centaur, on ConvNet and UCIHAR dataset dataset.
Correlation among Cost Metrics. To show the connection between cost metrics, we plot MAC, Latency, and Energy in Figure 8. We observe that both Latency and Energy have linear relationships with MAC. Besides, Latency is also linearly correlated with Energy, in such a way that these two entirely overlap when we appropriately scale the y-axes. We remark that Bandwidth also has a similar correlation. Such linear relationships are because of the assumption we made when computing Latency and Energy; however, the actual cost and the relationship among these metrics might not deviate much in practice. Based on such correlations, we can add Latency (or Energy) of model training as well as the communication Bandwidth to get the overall workload. The results shown in Figure 1 indicate that the workload still follows a similar pattern as the computation in model training, because the training workload significantly overweight the communication workload.
@@ 6.4. Data and Participation Heterogeneity
Centaur is robust to data heterogeneity. To create imbalanced non-IID data partitions among clients, we use LDA as defined in § 6.1. Next, we set different values for LDA-Alpha to manipulate the levels of non-IID data partitions. Figure 10 and Figure 10 show the results when LDA-Alpha are changing from $0.001$, to $0.1$, to $1$, and to $1000$. The smaller LDA-Alpha is, the less balanced the dataset is. We also annotate some examples of class distribution in Figure 10, demonstrating that LDA-Alpha=0.001=0.001 generates almost $1$ class per client, while LDA-Alpha=1000=1000 generates almost uniform classes per client. Results show that in all cases Centaur can reach higher test accuracy than both AP and UCD training. We remark that due to the retraining of the encoder, both Centaur and AP training can cause more fluctuations than UCD training; this is obvious when data distribution tends to be IID (e.g., when LDA-Alpha=1000=1000 in Figure 10).
Centaur scales well with higher total and participating clients. Here, we scale the FL training problem with the number of total clients from 10 to 1000. $10\%$, $20\%$, and $50\%$ of them are selected respectively as participating clients in each round. Figure 11 shows Centaur achieves higher test accuracy compared to both UCD training and AP training. Specifically, with a higher participation rate, all training methods tend to have better performance. However, with a much larger number of total clients (e.g., 1000), the test accuracy is reduced, because the same size of data is partitioned into more portions. Also, with such more fragmented data partitioning, Centaur has much accuracy gain when compared to UCD training. This experiment focuses on test accuracy and does not account for the orchestration costs of running Centaur in dynamic mobile environments. However, such scalability challenges are not unique to Centaur but are inherent to any FL framework. We also note that practical deployment at larger scales necessitates additional server-side coordination. Centaur is designed to be modular, enabling integration into existing FL stacks. Nevertheless, evaluating the end-to-end overhead of managing varying participation rates and asynchronous aggregation remains open.
@@ 6.5. Performance of Centaur under Mobility
Figure 11. The impact of participating vs. total clients on test accuracy for MobileNetv3 with CIFAR10.
@@
We study when the connection probability $\lambda$ changes based on the mobility model defined in §6.1. For a more readable visualization, we bucket $\lambda$ in three different ranges of $\lambda\in$ [0.1,0.4], $\lambda\in$ [0.4,0.7], and $\lambda\in$ [0.7,1.0] and report the corresponding accuracy when the UCDs are mobile. Figure 12 shows Centaur always has higher efficiency than standard UCD training while incurring a higher latency than direct AP training. Specifically, Centaur gains $\%, 5.645.64\%$, and $4.11\%$ higher accuracy with $51.73\%$, $60.48\%$, and $68.35\%$ lower cost both in terms of energy and latency across the different ranges of $\lambda$. Interestingly, we also observe a rise in accuracy with lower connectivity probability. The main reason is the design of our experimental setup: we realistically assume that with limited connectivity, UCDs devices can collect more data and perform data selection locally. Using this selected data, each device then trains the classifier for more epochs, which further improves its accuracy.
7. Discussion
While Centaur shows strong empirical performance and feasibility on constrained devices, some of our assumptions need further study. We outline key limitations, and opportunities for future work, below.
(1) Centaur assumes each UCD has a trusted, resourceful personal AP (e.g., a smartphone or router) that is always available. This simplifies privacy and connectivity but may not hold in transient or multi-user settings.
(2) We used RaspberryPi as a reproducible proxy for constrained devices, though it is more capable than many real UCDs. Results may thus overestimate feasibility; testing on truly constrained hardware is needed to validate our claims.
(3) Freezing the encoder and training only the classifier reduces on-device compute and memory but limits adaptability: if local data diverge from the pretrained one, classifier-only updates may be insufficient. Studying trade-offs in encoder vs. classifier update frequency remains open.
(4) Our on-device selection uses loss and gradient norm for importance sampling, but this remains heuristic without formal convergence or variance guarantees. It may also be sensitive to outliers, mislabeled, or out-of-distribution data, potentially biasing updates.
(5) Centaur reduces on-device workload but increases AP computation and network use. We did not quantify orchestration and scaling costs with many asynchronous users. Although fewer samples are sent than baselines, uplink and energy costs for constrained devices and overall communication efficiency need further analysis.
(6) We use FedAvg in our implementation, but Centaur can accommodate other FL aggregation strategies. Layer-wise FL methods that perform model training in a partitioned manner, such as FedMA wang2020federated, can also be integrated. Also, we evaluated Centaur on supervised tasks, yet a complementary approach can be suggested for unsupervised tasks.
@@ 8. Conclusions
We propose an efficient training method that leverages data selection to improve learning and reduce computational costs in multidevice federated learning. Our approach allows partition-based training and aggregation on resource-constrained devices, supported by resourceful companion devices. Our evaluations over various scenarios confirm that our proposed solution improves model accuracy while reducing system overhead.